Question 5e, f and g

The demand and supply curves for the market are given by:
Demand: Qd = 16000 – 24P
Supply: Qs = 2000 + 32P
A maximum price of $200 is proposed.

(e)   Does the imposition of the maximum price result in net welfare gains? Explain. (4 marks)
(f) Would you advise the implementation of the maximum price? Explain. (3 marks)
(g)   Suggest an alternate strategy. (2 marks)
Answer #1

(e)

Imposition of maximum price (a price ceiling) is binding if imposed below the equilibrium (free market price). At the lower price, quantity demanded rises but quantity supplied falls. Therefore, market quantity falls (to the level of quantity supplied). As a result, all intended beneficiaries from the price ceiling program are unable to buy the good at lower price, which, together with a decrease in market quantity, decreases net welfare gains by causing a net welfare (deadweight) loss.

(f)

In free market equilibrium, by equating Qd and Qs, we get P = $250.

When P = 200,

Qd = 16000 - (24 x 200) = 16000 - 4800 = 11200

Qs = 2000 + (32 x 200) = 2000 + 6400 = 8400

Market quantity = Qs = 8400.

When Q = 8400, from demand function,

8400 = 16000 - 24P

24P = 7600

P = $317

Deadweight loss (DWL) = (1/2) x $(317 - 200) x 8400 = 4200 x $117 = $491400

Since DWL > 0, I will not advise the implementation of the maximum price.

(g)

Social welfare is maximized when Qd = Qs.

16000 - 24P = 2000 + 32P

56P = 14000

P = $250

Q = 2000 + (32 x 250) = 2000 + 8000 = 10000

Therefore, this is the efficiency-maximizing price and quantity, and I would suggest non-intervention in the market by allowing free market.